OpenStudy (anonymous):

Let f (x) = 7x + sqrt(x + 50) . What is the form of the largest interval on which f has an inverse and the value(s) of the endpoint(s) find (f ^−1)' (0)

OpenStudy (anonymous):

@zepdrix @wio can you guys help?

OpenStudy (anonymous):

Hmmmm.

OpenStudy (anonymous):

I know how to find \(f^{-1}(0)\)

OpenStudy (anonymous):

Basically it's just: \[ 0= 7x + \sqrt{x + 50} \]

OpenStudy (anonymous):

Then multiply by \(7x-\sqrt{x+50}\).

OpenStudy (anonymous):

So \[ 0=(7x)^2-(\sqrt{x+50})^2 \]

OpenStudy (anonymous):

so is f^-1(0)=(7x)^2-x+50?

OpenStudy (anonymous):

That \(-\) should distribute to both terms.

OpenStudy (anonymous):

As far as an inverse goes... you can't invert it for \(x\) values where there is another \(x\) value has the same output.

OpenStudy (anonymous):

It's clear the lowest value of \(x\) where \(f(x)\) is defined is where: \[ x+50\geq 0 \]

OpenStudy (anonymous):

So \[ x\geq -50 \]

OpenStudy (anonymous):

your equation thing is not working. I dont understand what your writing.

OpenStudy (anonymous):

Oh wait... is it asking for derivative of inverse function?

OpenStudy (anonymous):

That makes so much more sense.

OpenStudy (anonymous):

yeah I think so

OpenStudy (anonymous):

also the interval

OpenStudy (anonymous):

If you find when the derivative is 0, then you find the max.

OpenStudy (anonymous):

max or min

OpenStudy (anonymous):

so the derivative is 7+(1/2)(x+50)^(-1/2)

OpenStudy (anonymous):

what is the interval?

OpenStudy (anonymous):

Find out when the derivative equalis 0 first.

OpenStudy (anonymous):

I got 1/196-50

OpenStudy (anonymous):

What does \(f(x)\) equal at that point?

OpenStudy (anonymous):

-349.89

OpenStudy (anonymous):

I think its supposed to be -(1/196)-50

OpenStudy (anonymous):

I think the interval would just be\[ \left[\frac{1}{196}-50,\infty\right) \]

OpenStudy (anonymous):

its not 1/196-50 but the interval is correct otherwise

OpenStudy (anonymous):

its [-50, infinity)